### v1

@@ 5.1 Galerkin boundary element discretization
(20) $C_{D}\left\lVert{(I_{h}^{-1}\circ S_{h})\boldsymbol{u}_{h}-S_{i\kappa}\boldsymbol{u}_{h}}\right\rVert_{\operatorname{\mathbf{X}}_{\Gamma}}$ $\leq\sup_{\begin{subarray}{c}\boldsymbol{w}\in\operatorname{\mathbf{X}}_{\Gamma}\\ \left\lVert{\boldsymbol{w}}\right\rVert_{\operatorname{\mathbf{X}}_{\Gamma}}=1\end{subarray}}\left|{\left\langle{\boldsymbol{w},(I_{h}^{-1}\circ S_{h})\overline{\boldsymbol{u}_{h}}-S_{i\kappa}\overline{\boldsymbol{u}_{h}}}\right\rangle_{\times,\Gamma}}\right|$
$\leq\widetilde{C}\left\lVert{\boldsymbol{u}_{h}}\right\rVert_{\operatorname{\mathbf{X}}_{\Gamma}}\sup_{\begin{subarray}{c}\boldsymbol{w}\in\operatorname{\mathbf{X}}_{\Gamma}\\ \left\lVert{\boldsymbol{w}}\right\rVert_{\operatorname{\mathbf{X}}_{\Gamma}}=1\end{subarray}}\inf_{\boldsymbol{w}_{h}\in\operatorname{\mathbf{U}}_{h}}\left\lVert{\boldsymbol{w}-\boldsymbol{w}_{h}}\right\rVert_{\operatorname{\mathbf{X}}_{\Gamma}}.$
Since the union of all lowest-order Raviart-Thomas boundary element spaces $\operatorname{\mathbf{U}}_{h}$, with $h>0$, is dense in $\operatorname{\mathbf{X}}_{\Gamma}$ (cf. [7, Corollary 5]), we have
$\varepsilon_{h}:=\widetilde{C}\,C_{D}^{-1}\sup_{\begin{subarray}{c}\boldsymbol{w}\in\operatorname{\mathbf{X}}_{\Gamma}\\ \left\lVert{\boldsymbol{w}}\right\rVert_{\operatorname{\mathbf{X}}_{\Gamma}}=1\end{subarray}}\inf_{\boldsymbol{w}_{h}\in\operatorname{\mathbf{U}}_{h}}\left\lVert{\boldsymbol{w}-\boldsymbol{w}_{h}}\right\rVert_{\operatorname{\mathbf{X}}_{\Gamma}}\to 0\qquad\quad\text{as}\quad h\to 0.$
Finally, let $T_{h}:=I_{h}^{-1}\circ S_{h}$. We arrive at the following estimate
$a(\boldsymbol{u}_{h},T_{h}\boldsymbol{u}_{h})$ $=a(\boldsymbol{u}_{h},S_{i\kappa}\boldsymbol{u}_{h})+a(\boldsymbol{u}_{h},(I_{h}^{-1}\circ S_{h})\boldsymbol{u}_{h}-S_{i\kappa}\boldsymbol{u}_{h})$
@@
$\geq C_{A}\left\lVert{\boldsymbol{u}_{h}}\right\rVert^{2}_{\operatorname{\mathbf{X}}_{\Gamma}}-\left\lVert{a}\right\rVert\left\lVert{\boldsymbol{u}_{h}}\right\rVert_{\operatorname{\mathbf{X}}_{\Gamma}}\left\lVert{(I_{h}^{-1}\circ S_{h})\boldsymbol{u}_{h}-S_{i\kappa}\boldsymbol{u}_{h}}\right\rVert_{\operatorname{\mathbf{X}}_{\Gamma}}$
$\geq\left({C_{A}-\varepsilon_{h}\left\lVert{a}\right\rVert}\right)\left\lVert{\boldsymbol{u}_{h}}\right\rVert^{2}_{\operatorname{\mathbf{X}}_{\Gamma}}.$
Choosing $h_{0}>0$ and $\widetilde{C}_{A}>0$ such that $C_{A}-\varepsilon_{h}\left\lVert{a}\right\rVert\geq\widetilde{C}_{A}$ for all $h<h_{0}$, we can complete the proof.
@@
Now, we shall prove the discrete inf-sup condition for the sesquilinear form $b:\operatorname{\mathbf{U}}_{h}\times\operatorname{\mathbf{V}}_{h}\to\mathbb{C}$, which is crucial in establishing the unique solvability of the discrete variational problem (16).
@@
Let Assumption 1 be satisfied. Then, there exist an $h_{0}>0$ and a constant $\widetilde{C}_{B}>0$ such that for all $h<h_{0}$ and for all $\boldsymbol{u}_{h}\in\operatorname{\mathbf{U}}_{h}$
(21) $\sup_{\boldsymbol{v}_{h}\in\operatorname{\mathbf{V}}_{h}\setminus\left\{{\boldsymbol{0}}\right\}}\dfrac{\left|{b(\boldsymbol{u}_{h},\boldsymbol{v}_{h})}\right|}{\left\lVert{\boldsymbol{v}_{h}}\right\rVert_{\operatorname{\mathbf{X}}_{\Gamma}}}\geq\widetilde{C}_{B}\left\lVert{\boldsymbol{u}_{h}}\right\rVert_{\operatorname{\mathbf{X}}_{\Gamma}}.$
Proof 5.4.
The discrete inf-sup condition (21) can be proven by following the approach in the proof of [25, Lemma 5.3].
The unique solvability of the discrete variational problem (16) and the asymptotic quasi-optimality of the discrete solutions immediately follow from Lemma 5.3, see [33, Theoreo 4.2.1].
@@
Let Assumption 1 be satisfied. Then, there exists an $h_{0}>0$ such that for all $h<h_{0}$, the discrete problem (16) has a unique solution $\boldsymbol{u}_{h}\in\operatorname{\mathbf{U}}_{h}$. In addition, the discrete solutions $\boldsymbol{u}_{h}$ converge to the solution $\boldsymbol{u}\in\operatorname{\mathbf{X}}_{\Gamma}$ of the problem (10) and satisfy the asymptotically quasi-optimal error estimate
(22) $\left\lVert{\boldsymbol{u}-\boldsymbol{u}_{h}}\right\rVert_{\operatorname{\mathbf{X}}_{\Gamma}}\leq C\inf_{\boldsymbol{w}_{h}\in\operatorname{\mathbf{U}}_{h}}\left\lVert{\boldsymbol{u}-\boldsymbol{w}_{h}}\right\rVert_{\operatorname{\mathbf{X}}_{\Gamma}},$
@@
It is worth mentioning that the quasi-optimal error estimate (22) alone does not imply the actual convergence rate of the Galerkin discretization scheme (16). When investigating the convergence rate of a Galerkin discretization scheme for Maxwell’s equations on polyhedral domains, we must take into account the singularities of the solutions, as shown by Costabel and Dauge in [17]. In particular, the error of best approximations by lowest-order Raviart-Thomas boundary elements was established in [22, Lemma 8.1], which is limited by a regularity exponent $s^{\ast}>0$ of the Laplace-Beltrami operator $\Delta_{\Gamma}$. The parameter $s^{\ast}$ depends only on the geometry of $\Gamma$ in neighborhoods of vertices. Moreover, there exist polyhedral vertices for which $s^{\ast}>0$ is arbitrarily small, as demonstrated in [5]. We refer the reader to [5, Theorem 8] and [22, Theorem 2.1] for more details on the regularity parameter $s^{\ast}$. Now, we are in the position to restate Lemma 8.1 in [22], which we exploit for our convergence analysis.
@@
$\inf_{\boldsymbol{w}_{h}\in\operatorname{\mathbf{U}}_{h}}\left\lVert{\boldsymbol{z}-\boldsymbol{w}_{h}}\right\rVert_{\operatorname{\mathbf{X}}_{\Gamma}}\leq Ch^{\min\left\{{\tfrac{3}{2}-\varepsilon,s+\tfrac{1}{2}-\varepsilon,1+s^{\ast},s+s^{\ast}}\right\}}\left\lVert{\boldsymbol{z}}\right\rVert_{\operatorname{\mathbf{H}}^{s}_{\times}(\textup{div}_{\Gamma},\Gamma)},$
@@
Let Assumption 1 be satisfied. In addition, we assume that the solution to (10) satisfies $\boldsymbol{u}\in\operatorname{\mathbf{H}}^{s}_{\times}(\textup{div}_{\Gamma},\Gamma),s>0$. Then, there exists an $h_{0}>0$ such that for all $h<h_{0}$ and for any $\varepsilon>0$, the unique solution $\boldsymbol{u}_{h}\in\operatorname{\mathbf{U}}_{h}$ to (16) satisfies
(23) $\left\lVert{\boldsymbol{u}-\boldsymbol{u}_{h}}\right\rVert_{\operatorname{\mathbf{X}}_{\Gamma}}\leq Ch^{\min\left\{{\tfrac{3}{2}-\varepsilon,s+\tfrac{1}{2}-\varepsilon,1+s^{\ast},s+s^{\ast}}\right\}}\left\lVert{\boldsymbol{u}}\right\rVert_{\operatorname{\mathbf{H}}^{s}_{\times}(\textup{div}_{\Gamma},\Gamma)},$
@@ 5.2 Matrix representation
Next, the Galerkin matrices of the sesquilinear forms $b(\cdot,\cdot)$ and $d(\cdot,\cdot)$ are defined by
${[\operatorname{\mathbf{B}}]}_{mn}:=b(\boldsymbol{u}_{n},\boldsymbol{v}_{m}),\qquad\qquad\quad{[\operatorname{\mathbf{D}}]}_{mn}:=d(\boldsymbol{u}_{n},\boldsymbol{v}_{m}),$
@@
Let Assumption 1 be satisfied. Then, there exists an $h_{0}>0$ such that for all $h<h_{0}$, the matrix equation (25) has a unique solution $\widehat{\boldsymbol{u}}_{h}\in\mathbb{C}^{N}$. In addition, the corresponding $\boldsymbol{u}_{h}\in\operatorname{\mathbf{U}}_{h}$ defined by (24) is a quasi-optimal approximation to the solution $\boldsymbol{u}$ of (10) in the sense of (22).
@@
The following result shows that the problem (25) is well-conditioned regardless of the boundary mesh resolution.
@@
Let Assumption 1 be satisfied. Then, there exist an $h_{0}>0$ and a constant $C>0$ such that for all $h<h_{0}$
@@
Proof 5.10.
It holds that
$\left|{\lambda_{\max}(\operatorname{\mathbf{D}}^{-1}\operatorname{\mathbf{B}})}\right|$ $\leq\left\lVert{\operatorname{\mathbf{D}}^{-1}\operatorname{\mathbf{B}}}\right\rVert\leq\widetilde{C}_{D}^{-1}\left\lVert{b}\right\rVert,$
$\left|{\lambda_{\max}(\operatorname{\mathbf{B}}^{-1}\operatorname{\mathbf{D}})}\right|$ $\leq\left\lVert{\operatorname{\mathbf{B}}^{-1}\operatorname{\mathbf{D}}}\right\rVert\leq\widetilde{C}_{B}^{-1}\left\lVert{d}\right\rVert.$

### v2

@@ 5.1 Galerkin boundary element discretization
$\quad-\sup_{\boldsymbol{u}_{h}\in\operatorname{\mathbf{U}}_{h}}\sup_{\boldsymbol{v}_{h}\in\operatorname{\mathbf{V}}_{h}}\dfrac{\left|{\left\langle{C_{i\kappa^{\prime}}\boldsymbol{u}_{h},\overline{\boldsymbol{v}}_{h}}\right\rangle_{\times,\Gamma}}\right|}{\left\lVert{\boldsymbol{u}_{h}}\right\rVert_{\operatorname{\mathbf{X}}_{\kappa^{\prime}}}\left\lVert{S^{-1}_{i\kappa^{\prime}}\boldsymbol{v}_{h}}\right\rVert_{\operatorname{\mathbf{X}}_{\kappa^{\prime}}}}$
@@
$\geq\dfrac{\alpha}{2}-\sup_{\boldsymbol{u}_{h}\in\operatorname{\mathbf{U}}_{h}}\dfrac{\left\lVert{C_{i\kappa^{\prime}}\boldsymbol{u}_{h}}\right\rVert_{\operatorname{\mathbf{X}}_{\kappa^{\prime}}}}{\left\lVert{\boldsymbol{u}_{h}}\right\rVert_{\operatorname{\mathbf{X}}_{\kappa^{\prime}}}}\geq\dfrac{\alpha}{2}-\beta.$
The constant $\tfrac{\alpha}{2}-\beta$ is only dependent of $\kappa^{\prime}$ and $\Gamma$, but independent of $h$. Moreover, it can be non-positive. In order to ensure that the inf-sup condition (21) is satisfied with a positive constant, we make the following assumption. {assumption}[Stability condition] There is a real number $\kappa^{\prime}_{0}>0$ such that
(22) $\alpha(\kappa^{\prime}_{0},\Gamma)>2\beta(\kappa^{\prime}_{0},\Gamma),$
where $\alpha$ and $\beta$ are defined by (19) and (10), respectively.
@@
Now, we shall show the discrete inf-sup condition for the sesquilinear form $b:\operatorname{\mathbf{U}}_{h}\times\operatorname{\mathbf{V}}_{h}\to\mathbb{C}$, which is crucial in establishing the unique solvability of the discrete variational problem (20). The next result can be proven by following the hints in [22, Exercise 26.5], with the approximability of $\operatorname{\mathbf{U}}_{h}$ and $\operatorname{\mathbf{V}}_{h}$ in $\operatorname{\mathbf{H}}^{-1/2}_{\times}(\textup{div}_{\Gamma},\Gamma)$ stated in [8, Corollary 5] and [3], respectively. We leave the complete proof for the reader.
@@
Let $\kappa$ and $\kappa^{\prime}_{0}$ fulfill Assumptions 1 and 5.2, respectively. Then, there exist an $h_{0}>0$ and a constant $\gamma>0$ such that for all $h<h_{0}$
(23) $\inf_{\boldsymbol{u}_{h}\in\operatorname{\mathbf{U}}_{h}}\sup_{\boldsymbol{v}_{h}\in\operatorname{\mathbf{V}}_{h}}\dfrac{\left|{b(\boldsymbol{u}_{h},\boldsymbol{v}_{h})}\right|}{\left\lVert{\boldsymbol{u}_{h}}\right\rVert_{\operatorname{\mathbf{X}}_{\kappa_{0}^{\prime}}}\left\lVert{S^{-1}_{i\kappa_{0}^{\prime}}\boldsymbol{v}_{h}}\right\rVert_{\operatorname{\mathbf{X}}_{\kappa_{0}^{\prime}}}}\geq\gamma.$
The unique solvability of the discrete variational problem (20) and the asymptotic quasi-optimality of the discrete solutions immediately follow from Lemma 5.3, see [37, Theorem 4.2.1].
@@
Let Assumptions 1 and 5.2 be satisfied. Then, there exists an $h_{0}>0$ such that for all $h<h_{0}$, the discrete problem (20) has a unique solution $\boldsymbol{u}_{h}\in\operatorname{\mathbf{U}}_{h}$. In addition, the discrete solutions $\boldsymbol{u}_{h}$ converge to the solution $\boldsymbol{u}\in\operatorname{\mathbf{H}}^{-1/2}_{\times}(\textup{div}_{\Gamma},\Gamma)$ of the problem (15) and satisfy the asymptotically quasi-optimal error estimate
(24) $\left\lVert{\boldsymbol{u}-\boldsymbol{u}_{h}}\right\rVert_{\operatorname{\mathbf{H}}^{-1/2}_{\times}(\textup{div}_{\Gamma},\Gamma)}\leq C\inf_{\boldsymbol{w}_{h}\in\operatorname{\mathbf{U}}_{h}}\left\lVert{\boldsymbol{u}-\boldsymbol{w}_{h}}\right\rVert_{\operatorname{\mathbf{H}}^{-1/2}_{\times}(\textup{div}_{\Gamma},\Gamma)},$
@@
It is worth mentioning that the quasi-optimal error estimate (24) alone does not imply the actual convergence rate of the Galerkin discretization scheme (20). When investigating the convergence rate of a Galerkin discretization scheme for Maxwell’s equations on polyhedral domains, we must take into account the singularities of the solutions, as shown by Costabel and Dauge in [19]. In particular, the error of best approximations by lowest-order Raviart-Thomas boundary elements was established in [26, Lemma 8.1], which is limited by a regularity exponent $s^{\ast}>0$ of the Laplace-Beltrami operator $\Delta_{\Gamma}$. The parameter $s^{\ast}$ depends only on the geometry of $\Gamma$ in neighborhoods of vertices. Moreover, there exist polyhedral vertices for which $s^{\ast}>0$ is arbitrarily small, as demonstrated in [6]. We refer the reader to [6, Theorem 8] and [26, Theorem 2.1] for more details on the regularity parameter $s^{\ast}$. Now, we are in the position to restate Lemma 8.1 in [26], which we exploit for our convergence analysis.
@@
$\inf_{\boldsymbol{w}_{h}\in\operatorname{\mathbf{U}}_{h}}\left\lVert{\boldsymbol{z}-\boldsymbol{w}_{h}}\right\rVert_{\operatorname{\mathbf{H}}^{-1/2}_{\times}(\textup{div}_{\Gamma},\Gamma)}\leq Ch^{\min\left\{{\tfrac{3}{2}-\varepsilon,s+\tfrac{1}{2}-\varepsilon,1+s^{\ast},s+s^{\ast}}\right\}}\left\lVert{\boldsymbol{z}}\right\rVert_{\operatorname{\mathbf{H}}^{s}_{\times}(\textup{div}_{\Gamma},\Gamma)},$
@@
Let Assumptions 1 and 5.2 be satisfied. In addition, we assume that the solution to (15) satisfies $\boldsymbol{u}\in\operatorname{\mathbf{H}}^{s}_{\times}(\textup{div}_{\Gamma},\Gamma),s>0$. Then, there exists an $h_{0}>0$ such that for all $h<h_{0}$ and for any $\varepsilon>0$, the unique solution $\boldsymbol{u}_{h}\in\operatorname{\mathbf{U}}_{h}$ to (20) satisfies
(25) $\left\lVert{\boldsymbol{u}-\boldsymbol{u}_{h}}\right\rVert_{\operatorname{\mathbf{H}}^{-1/2}_{\times}(\textup{div}_{\Gamma},\Gamma)}\leq Ch^{\min\left\{{\tfrac{3}{2}-\varepsilon,s+\tfrac{1}{2}-\varepsilon,1+s^{\ast},s+s^{\ast}}\right\}}\left\lVert{\boldsymbol{u}}\right\rVert_{\operatorname{\mathbf{H}}^{s}_{\times}(\textup{div}_{\Gamma},\Gamma)},$
@@ 5.2 Matrix representation
Next, the Galerkin matrices of the sesquilinear form $b(\cdot,\cdot)$ and the pairing ,×$\left\langle{\cdot,\cdot}\right\rangle_{\times,\Gamma}$ are respectively defined by
${[\operatorname{\mathbf{B}}]}_{mn}:=b(\boldsymbol{u}_{n},\boldsymbol{v}_{m}),\qquad\qquad\quad{[\operatorname{\mathbf{D}}]}_{mn}:=\left\langle{\boldsymbol{u}_{n},\overline{\boldsymbol{v}_{m}}}\right\rangle_{\times,\Gamma},$
@@
Let Assumptions 1 and 5.2 be satisfied. Then, there exists an $h_{0}>0$ such that for all $h<h_{0}$, the matrix equation (27) has a unique solution $\widehat{\boldsymbol{u}}_{h}\in\mathbb{C}^{N}$. In addition, the corresponding $\boldsymbol{u}_{h}\in\operatorname{\mathbf{U}}_{h}$ defined by (26) is a quasi-optimal approximation to the solution $\boldsymbol{u}$ of (15) in the sense of (24).
@@
The following result shows that the problem (27) is well-conditioned regardless of the boundary mesh refinement.
@@
Let Assumptions 1 and 5.2 be satisfied. Then, there exist an $h_{0}>0$ and a constant $C>0$ such that for all $h<h_{0}$
@@
Proof 5.9.
We follow the proof of [25, Theorem 2.1]. Let the subspaces $\operatorname{\mathbf{U}}_{h}$ and $\operatorname{\mathbf{V}}_{h}$ be equipped with the norm $\left\lVert{\cdot}\right\rVert_{\operatorname{\mathbf{X}}_{\kappa^{\prime}_{0}}}$, where $\kappa^{\prime}_{0}$ fulfills Assumption 5.2. We denote by $B_{h}:\operatorname{\mathbf{U}}_{h}\to\operatorname{\mathbf{V}}_{h}^{\prime}$ and $D_{h}:\operatorname{\mathbf{U}}_{h}\to\operatorname{\mathbf{V}}_{h}^{\prime}$ the bounded linear operators associated with the sesquilinear form $b(\cdot,\cdot)$ and the pairing ,×$\left\langle{\cdot,\cdot}\right\rangle_{\times,\Gamma}$, respectively. The following estimates hold true due to (16), (19) and (23)
$\left\lVert{D_{h}}\right\rVert_{\operatorname{\mathbf{U}}_{h}\to\operatorname{\mathbf{V}}_{h}^{\prime}}\leq 1,$ $\left\lVert{D_{h}^{-1}}\right\rVert_{\operatorname{\mathbf{V}}_{h}^{\prime}\to\operatorname{\mathbf{U}}_{h}}\leq\alpha^{-1},$
$\left\lVert{B_{h}}\right\rVert_{\operatorname{\mathbf{U}}_{h}\to\operatorname{\mathbf{V}}_{h}^{\prime}}\leq\left\lVert{b}\right\rVert,$ $\left\lVert{B_{h}^{-1}}\right\rVert_{\operatorname{\mathbf{V}}_{h}^{\prime}\to\operatorname{\mathbf{U}}_{h}}\leq\gamma^{-1}.$
Next, let $\mathbb{C}^{N}$ be equipped with the norm $\left\lVert{\cdot}\right\rVert_{\operatorname{\mathbf{X}}_{\kappa^{\prime}_{0}}}$ via the coefficient isomorphism with respect to the basis $\{\boldsymbol{u}_{1},\boldsymbol{u}_{2},\ldots,\boldsymbol{u}_{N}\}$. Then, there hold that
$\left|{\lambda_{\max}(\operatorname{\mathbf{D}}^{-1}\operatorname{\mathbf{B}})}\right|$ $\leq\left\lVert{\operatorname{\mathbf{D}}^{-1}\operatorname{\mathbf{B}}}\right\rVert\leq\alpha^{-1}\left\lVert{b}\right\rVert,$
$\left|{\lambda_{\max}(\operatorname{\mathbf{B}}^{-1}\operatorname{\mathbf{D}})}\right|$ $\leq\left\lVert{\operatorname{\mathbf{B}}^{-1}\operatorname{\mathbf{D}}}\right\rVert\leq\gamma^{-1}.$